OpenStudy (love.kat123):

In 2005, a Nevada silver mine exported 200 tons of silver. In 2008, as the silver supply became exhausted, it only exported 25 tons of silver. What is the rate of change of the mine’s silver output? A. -175/3 B. -1/50 C. 175/3 D. 75

OpenStudy (love.kat123):

@texaschic101

OpenStudy (texaschic101):

you might need a second opinion...but I think it is : this is a slope problem slope = (y2 - y1) / (x2 - x1) (2005,200)...x1 = 2005 and y1 = 200 (2008,25)...x2 = 2008 and y2 = 25 now we sub slope = (25 - 200) / (2008 - 2005) slope = -175/3 rate of change = -175/3

OpenStudy (love.kat123):

thank u

OpenStudy (texaschic101):

don't thank me yet....not until I know if I did it correctly

OpenStudy (love.kat123):

well i already did so haha and u were right